"the length of a rectangle is 5 cm. more than the width, and the area is 84 cm. find the dimentions."
Answer by unlockmath(1688)

Hello,
First off, we know to find the area of a rectangle is width times height.
Let's say the width is represented by x and the length will be shown by x + 5.
Here's the equation:
(x)(x + 5)= 84 cm squared. multiply x and (x + 5) gives us:
x^2 + 5x = 84 (Note:^2 means to the second power or squared.)
x^2 + 5x - 84 = 0 Subtract 84 on both sides and then we can factor
(x + 12)(x - 7) = 0 We have 2 answers x = -12 and x = 7
Obviously -12 doesn't make sense for a measure so
x = 7 is the answer.
Width = 7 cm
Length = 12 cm Multiply these two numbers and the area is 84 cm sq
RJ Toftness
www.math-unlock.com